Please help me solve this problem.


A family, traveling along an interstate highway, at a legal speed, noted with suprise that their car's odometer reading was 45954, palindromic number (one which reads the same forward and backwards.) The family thought that such a palindromic mileage reading was unlikely to occur again in the near future. However, two hours later their car's odometer displayed yet another palindrome. Determine, in miles per hour, the car's speed during that two-hour period. Explain your answer in details.

_________________________________ miles per hour

Answer by ptaylor(2198)

Distance(d) equals Rate(r) times Time(t) or d=rt; r=d/t and t=d/r
By inspection, we see that the next palindromic number would have to be 46064,so:
46064 minus 45954=110 mi
r=d/t=110/2=55 mph------------------------ans